The average flight from point A to B takes 40 minutes with a standard deviation of 5 minutes. What is the probability a random sample of 4 flights will take between 38 and 43 minutes?

Answer by stanbon(75887)

The average flight from point A to B takes 40 minutes with a standard deviation of 5 minutes. What is the probability a random sample of 4 flights will take between 38 and 43 minutes?
---------------------------------------------------------
z(38) = (38-40)/[5/[sqrt(4)] = -0.8
----------
z(43) = (43-40)/[5/sqrt(4)] = 1.2
---------------------------
P(38 < x < 43) = P(-0.8 < z < 1.2) = 0.6731
==================================================
Cheers,
Stan H.
------------